Sum of Squared Error (SSE) measures the dispersion between actual and predicted values of the dependent variable.

true or false
Answer #1

TRUE.

[ explanation:-

Sum of squares (SS) is used to identify the dispersion of data.

in regression there are three sum of squares : SSE, SST, SSR

SSE is defined as the sum of squared errors..i.e, sum of the squared deviations between the observed and predicted values.

SSE=\sum e_{i}^2=\sum \left ( y_{i}-\hat{y_{i}} \right )^2 ]

*** if you have any doubt regarding the problem ,please write it in the comment box...if satisfied,please UPVOTE.